twice a number is 12 less than 6 times the opposite of the number: twice a number is 12 less than 6 times the opposite of the number
Answer by oscargut(667)

twice a number is 12 less than 6 times the opposite of the number
if we denote x as the number
2x=-6x-12
8x=-12
x=-12/8= -3/2
the number is -3/2
Question 146111: twice a number is 12 less than 6 times the opposite of the number: twice a number is 12 less than 6 times the opposite of the number
Answer by ankor@dixie-net.com(4538)

Let x = "a number"
then
-x = "opposite of the number'
:
Just write what it says (the word "is" usually means "=":
twice a number is 12 less than 6 times the opposite of the number
2x = 6(-x) - 12
2x = -6x - 12
2x + 6x = -12
8x = 12
x= (-12)/8
x = -1.5; (the opposite would be +1.5)
:
:
Check solution in original statement
2(-1.5) = 6(+1.5) - 12
-3 = +9 - 12